find the general formula for (-1/4, 2/9, -3/16, 4/25,...)

Answer by Theo(13342)

looks like the formula is x divided by (x+1)^2 * (-1)^2

when x = 1, the term is 1/2^2 * (-1)^1 = -1/4

when x = 2, the term is 2/3^2 * (-1)^2 = 2/9

whehn x = 3, the term is 3/4^2 * (-1)^3 = -3/16

when x = 4, the term is 4/5^2 * (-1)^4 = 4/25